Grade 12Physical Chemistry

Which is isoelectronic with no3 1.n22.co3.n2o4.no2+

Raghavendra
9 Years agoGrade 12

1 Answer

Sneh Agrawal
9 Years ago
No. of electrons in:NO3=7+24=31_____N2=7*2=14____CO=6+8=14_____N2O=7*2 + 8=22_____NO2+=7+(8*2)-1=22.From all the calculated number of electrons, we can conclude that N2 is isoelectronic with CO and N2O is isoelectronic with NO2+. We can also conclude that none of them is isoelectronic with NO3.______HOPE IT HELPED. BEST OF LUCK:)